# 4cm4cm3cm3cm

### v1

@@ Abstract.
M. Berezovik is partially supported by the ISF grant 938/22, M. Bialy is partially supported by ISF grant 974/24.
@@ 1. Introduction
[…]
@@ 2. Preliminaries
In convex geometry, it is well known that the body $X^{\circ}$ has a $C^{1}$-smooth boundary if and only if $X$ is strictly convex. In our case, we have $X=X^{\omega}=JX^{\circ}$. Therefore, $X$ and $X^{\circ}$ coincide up to rotation. Hence, $X$ is strictly convex if and only if $X$ has $C^{1}$-smooth boundary ∎
@@
Observe that the characteristic line of $\partial X$ at $x$ (which is defined as $\ker\omega|_{T_{x}\partial X}\subseteq T_{x}\partial X$) is parallel to $f(x)$ (this fact was noticed before in [11, Proof of Lemma 3.6]). Indeed, from the definition of $X^{\omega}$, we have $\omega(y,f(x))\leq 1$ for every $y\in X$. Hence, the hyperplane $\omega(\cdot,f(x))=1$ is a tangent hyperplane of $\partial X$ at the point $x$. Then it is easy to see that $f(x)$ is actually parallel to the characteristic line in this hyperplane. Moreover, one can show that $f$ is just a restriction of a Hamiltonian vector field generated by the Hamiltonian function $H=\|\cdot\|_{X}$ on $\partial X$. Here $\|\cdot\|_{X}$ is a norm on $\mathbb{R}^{2n}$ whose unit ball is $X$. Therefore, if $X$ has a $C^{k}$-smooth boundary $\partial X$, for $k\geq 1$, then $f$ is $C^{k-1}$-smooth map, as a map from $\partial X$ to $\mathbb{R}^{2n}$.
@@
Observe that $f\colon\partial X\to\partial X$ is a $C^{1}$-smooth map and $f^{2}=-\operatorname{id}_{\partial X}$. Therefore, $f$ is $C^{1}$-diffeomorphism of $\partial X$.
Arguing by contradiction, assume that there exists a point $x^{*}\in\partial X$ and a non-zero vector $\xi\in T_{x^{*}}\partial X$ such that $\nabla_{\xi}n=0$, where $n(x)$ is the unit outward-pointing normal vector field on $\partial X$.
We write $f(x)=\lambda(x)\cdot Jn(x)$ for a positive $C^{1}$-smooth function $\lambda$ on $\partial X$. Define the function $g\colon\partial X\to\mathbb{R}$ by $g(x)=\omega(x^{*},f(x))$. The function $g$ attains its maximum at the point $x^{*}$, giving us $\xi(g)=0$. At the same time, we have
@@
Therefore $\xi(\lambda)=0$. Moreover,
$\nabla_{\xi}(f)=\xi(\lambda)\cdot Jn(x^{*})+\lambda(x^{*})\cdot J\nabla_{\xi}n,$
hence $f_{*}(\xi)=0$. But this contradicts $f$ being a diffeomorphism. Therefore, $\partial X$ has positive curvature at every point.
@@
Remark 3.1.
The map $T$ is only continuous in general, if we assume that $\partial X$ is $C^{k}$-smooth and has positive curvature, then $T$ is $C^{k-1}$-smooth symplectomorphism.
@@
Let us show that $Tz_{4}=z_{1}$. It was observed before that the vector $f(x)$ gives the characteristic direction at the point $x$. Therefore, the line $z_{4}z_{1}$ coincides with the characteristic line at the point $x$ and the point $x$ divides the segment $[z_{4},z_{1}]$ in half. From the definition of the map $T$ we have $Tz_{4}=z_{1}$.
Note that at the point $y=f(x)\in\partial X$, the characteristic direction is given by vector $f(y)=f(f(x))=-x$. Therefore, $Tz_{1}=z_{2}$. The other cases follow from central symmetry. ∎
@@
Consider some $x\in\partial X$. Take point $z$ on the corresponding centrally symmetric $4$-periodic trajectory such that $x=(z+Tz)/2$, and consider point $y=(Tz+T^{2}z)/2\in\partial X$. From the central symmetry of the trajectory we have $y=(Tz-z)/2=x-z$, i.e. $y$ is parallel to the characteristic line at $x$, so now $y$ may not coincide with $f(x)$ but still co-directed, so $y=f(x)/\|f(x)\|_{X}$. For the same reason $x$ is parallel to the characteristic line at $y$.
@@ 4. Invariant hypersurface
From Theorem 3.2 we have that for symplectically self-polar convex body $X$ with a $C^{1}$-smooth boundary the set $Y=\{x+f(x):x\in\partial X\}$ is invariant under an outer billiard map. Indeed,
@@
In this section, we discuss the geometric properties of $Y$.
@@
The set $Y$ is a topological sphere and $X$ lies in the bounded component of $\mathbb{R}^{2n}\setminus Y$. If $X$ has $C^{k}$-smooth boundary for $k\geq 2$, then $Y$ is a $C^{k-1}$-embedded submanifold.
@@
To prove that $Y$ is a boundary of a star-shaped body, we have to prove that every ray of the form $\{tv:t\in[0,+\infty)\}$ intersects $Y$ exactly at one point for every $v\in\mathbb{R}^{2n}\setminus\{0\}$. From Lemma 4.1, we know that for every $v\in\mathbb{R}^{2n}\setminus\{0\}$ the corresponding ray intersects $Y$ at least at one point. Assume now that $y+f(y)=t(x+f(x))$ for some $x,y\in\partial X$ and $t\geq 1$. Consider symplectic subspace $V=\langle x,f(x)\rangle$, then $\mathbb{R}^{2n}=V\oplus V^{\omega}$, where $V^{\omega}=\{u\in\mathbb{R}^{2n}:\,\omega(u,w)=0\ \forall w\in V\}$. Let $y=y^{\parallel}+y^{\perp}$ and $f(y)=f(y)^{\parallel}+f(y)^{\perp}$ with respect to this decomposition. From $y+f(y)\in V$ we have $y^{\perp}+f(y)^{\perp}=0$. Since $y^{\parallel},f(y)^{\parallel}\in V$, we can write them as
@@
Now, let $\partial X$ be $C^{2}$-smooth. Notice that the tangent space of $Y$ at the point $x+f(x)$ for $x\in\partial X$ can be described as $\{\xi+\nabla_{\xi}f:\xi\in T_{x}\partial X\}$. Consider symplectic subspaces $V=\langle x,f(x)\rangle$, $V^{\omega}$. Remember that $f(x)$ and $x$ give characteristic directions in $T_{x}\partial X$ and $T_{f(x)}\partial X$ respectively. Therefore, by the definition of characteristic direction we have
@@
Every line of the form $\{x+tf(x):t\in\mathbb{R}\}$ intersects $Y$ exactly at two points $x+f(x)$ and $x-f(x)$ for every $x\in\partial X$. If $\partial X$ is $C^{2}$-smooth, then every such line is transversal to $Y$.
@@
Now, let $\partial X$ be $C^{2}$-smooth. Arguing by contradiction, assume that the line $\{x+tf(x):t\in\mathbb{R}\}$ intersects $Y$ not transversally at the point $x+f(x)$ for $x\in\partial X$ (case $x-f(x)$ is analogous). Following the proof of Theorem 4.2, there exists $\xi\in T_{x}\partial X$ such that
@@ 6. Discussion and open questions
In dimension two, for bodies with a $C^{2}$-smooth boundary, the answer is negative. Indeed, consider a symplectically self-polar convex body $X\subset\mathbb{R}^{2}$ and the corresponding set $Y$. Then $\partial X$ can be recovered from $Y$ by the so-called area construction (the outer billiard analog of the string construction [16]) as follows. Let $A$ be the area of the region bounded by $Y$. For any point $y$, consider the segment $[y,z(y)],z(y)\in Y$ cutting from $Y$ a segment of a constant area equal to $\frac{A-4}{4}$ (Figure 1, gray area). Then,
@@
Figure 1. Area construction applied to $Y$ with the area parameter $\frac{A-4}{4}$
@@
In [7], analytic examples of Radon curves were found via a Lamé equation. Interesingly, are there analytic examples of symplectically self-polar bodies in higher dimensions?

### v2

# Outer billiards of symplectically self-polar convex bodies

@@ Abstract.
2020 Mathematics Subject Classification:
37C05, 52A21, 53A07, 53D99
Mark Berezovik was partially supported by the ISF grant 938/22 and partially supported by the ISF-NSFC grant 3231/23.
Misha Bialy was partially supported by ISF grant 974/24.
@@ 1. Introduction
where $\langle\cdot,\cdot\rangle$ denotes the standard inner product in $\mathbb{R}^{2n}$. The relation between $X^{\omega}$ and $X^{\circ}$ is given by $X^{\omega}=JX^{\circ}$, where $J$ is the multiplication by $\sqrt{-1}$ under the standard identification of $\mathbb{C}^{n}\cong\mathbb{R}^{2n}$.
@@
Acknowledgments.
The authors thank Roman Karasev and the anonymous referee for useful remarks and comments.
@@ 2. Preliminaries
In convex geometry, it is well known that the body $X^{\circ}$ has a $C^{1}$-smooth boundary if and only if $X$ is strictly convex. In our case, we have $X=X^{\omega}=JX^{\circ}$. Therefore, $X$ and $X^{\circ}$ coincide up to rotation. Hence, $X$ is strictly convex if and only if $X$ has a $C^{1}$-smooth boundary. ∎
@@
Observe that the characteristic line of $\partial X$ at $x$ (which is defined as $\ker\omega|_{T_{x}\partial X}\subseteq T_{x}\partial X$) is parallel to $f(x)$ (this fact was noticed before in [11, Proof of Lemma 3.6]). Indeed, from the definition of $X^{\omega}$, we have $\omega(y,f(x))\leq 1$ for every $y\in X$. Hence, the hyperplane $\omega(\cdot,f(x))=1$ is a tangent hyperplane of $\partial X$ at the point $x$. Then it is easy to see that $f(x)$ is actually parallel to the characteristic line in this hyperplane. Moreover, one can show that $f$ is just a restriction of a Hamiltonian vector field generated by the Hamiltonian function $H=\|\cdot\|_{X}$ on $\partial X$. Here $\|\cdot\|_{X}$ is a norm on $\mathbb{R}^{2n}$ whose unit ball is $X$. Precisely, for $x\in\mathbb{R}^{2n}$,
$\|x\|_{X}=\inf\{\lambda>0:\lambda^{-1}x\in X\}.$
Therefore, if $X$ has a $C^{k}$-smooth boundary $\partial X$, for $k\geq 1$, then $f$ is $C^{k-1}$-smooth map, as a map from $\partial X$ to $\mathbb{R}^{2n}$.
For a symplectically self-polar convex body $X\subset\mathbb{R}^{2n}$ the map $f$ acts from $\partial X$ to itself. Therefore, we can consider its composition with itself. Although the next statement is simple, we state it explicitly, since it is used throughout the whole paper.
Lemma 2.2.
Let $X\subset\mathbb{R}^{2n}$ be a symplectically self-polar convex body with a $C^{1}$-smooth boundary. Then $f\circ f=-\operatorname{id}_{\partial X}$.
Proof.
Let $x\in\partial X$, then by the definition of $f$ we have $\omega(x,f(x))=1$. Denote $y=f(x)\in\partial X$, then $\omega(y,-x)=\omega(x,f(x))=1$. Since $X$ is centrally symmetric, it follows that $x\in\partial X=\partial X^{\omega}$. By the definition of $f$ this means that f(y)=xf(y)=-x. ∎
@@
Observe that $f\colon\partial X\to\partial X$ is a $C^{1}$-smooth map and $f^{2}=-\operatorname{id}_{\partial X}$. Therefore, $f$ is a $C^{1}$-diffeomorphism of $\partial X$.
Assume, for the sake of contradiction, that there exists a point $x^{*}\in\partial X$ and a non-zero vector $\xi\in T_{x^{*}}\partial X$ such that $\nabla_{\xi}n=0$, where $n(x)$ is the unit outward-pointing normal vector field on $\partial X$.
We write $f(x)=\lambda(x)\cdot Jn(x)$ for a positive $C^{1}$-smooth function $\lambda$ on $\partial X$. Define the function $g\colon\partial X\to\mathbb{R}$ by $g(x)=\omega(x^{*},f(x))$. The function $g$ attains its maximum at the point $x^{*}$, giving us $\xi(g)=0$, where $\xi(g)$ is the derivative of $g$ with respect to $\xi$. At the same time, we have
@@
Therefore, $\xi(\lambda)=0$. Hence $f_{*}(\xi)=0$, since
$f_{*}(\xi)=\nabla_{\xi}(f)=\xi(\lambda)\cdot Jn(x^{*})+\lambda(x^{*})\cdot J\nabla_{\xi}n,$
where $f_{*}$ is the differential of $f$. This contradicts $f$ being a diffeomorphism. Therefore, $\partial X$ has positive curvature at every point.
@@
Figure 1. The symplectic outer billiard map in the plane
Reversing the orientation of the characteristic line bundle, one gets the inverse map $T^{-1}$. Therefore, the map $T$ is a bijection of $\mathbb{R}^{2n}\setminus X$.
The map $T$ is only continuous in general, if we assume that $\partial X$ is $C^{k}$-smooth and has positive curvature, then $T$ is a $C^{k-1}$-smooth symplectomorphism.
@@
Let us show that $Tz_{4}=z_{1}$. It was observed before that the vector $f(x)$ gives the characteristic direction at the point $x$. Therefore, the line $z_{4}z_{1}$ coincides with the characteristic line at the point $x$ and the point $x$ divides the segment $z_{4},z_{1}]$ in half. From the definition of the map $T$ we have $Tz_{4}=z_{1}$.
By Lemma 2.2 the characteristic direction at the point $y=f(x)\in\partial X$ is given by the vector f(y)=f(f(x))=xf(y)=f(f(x))=-x. Therefore, $Tz_{1}=z_{2}$. The other cases follow from central symmetry. ∎
@@
Consider some $x\in\partial X$. Take point $z$ on the corresponding centrally symmetric $4$-periodic trajectory such that $x=(z+Tz)/2$, and consider point $y=(Tz+T^{2}z)/2\in\partial X$. From the central symmetry of the trajectory we have y=(Tzz)/2=xzy=(Tz-z)/2=x-z, i.e. $y$ is parallel to the characteristic line at $x$, so now $y$ may not coincide with $f(x)$ but still co-directed, so $y=f(x)/\|f(x)\|_{X}$. For the same reason $x$ is parallel to the characteristic line at $y$.
@@ 4. Invariant hypersurface
Throughout this section we assume $X\subset\mathbb{R}^{2n}$ to be a symplectically self-polar convex body. Theorem 3.1 and Lemma 2.2 imply that for such an $X$ with a $C^{1}$-smooth boundary the set $Y=\{x+f(x):x\in\partial X\}$ is invariant under an outer billiard map. Indeed,
@@
In this section we discuss the geometric properties of $Y$.
@@
The set $Y$ is a topological sphere, and $X$ lies in the bounded component of $\mathbb{R}^{2n}\setminus Y$. If $X$ has $C^{k}$-smooth boundary for $k\geq 2$, then $Y$ is a $C^{k-1}$-embedded submanifold.
@@
To prove that $Y$ is a boundary of a star-shaped body, we have to prove that every ray of the form $\{tv:t\in[0,+\infty)\}$ intersects $Y$ exactly at one point for every $v\in\mathbb{R}^{2n}\setminus\{0\}$. From Lemma 4.1, we know that for every $v\in\mathbb{R}^{2n}\setminus\{0\}$ the corresponding ray intersects $Y$ at least at one point. Assume now that $y+f(y)=t(x+f(x))$ for some $x,y\in\partial X$ and $t\geq 1$. Consider the symplectic subspace $V=\mathop{\text{span}}\{x,f(x)\}$, then $\mathbb{R}^{2n}=V\oplus V^{\omega}$, where $V^{\omega}=\{u\in\mathbb{R}^{2n}:\,\omega(u,w)=0\ \forall w\in V\}$. Let $y=y^{\parallel}+y^{\perp}$ and $f(y)=f(y)^{\parallel}+f(y)^{\perp}$ with respect to this decomposition. From $y+f(y)\in V$ we have $y^{\perp}+f(y)^{\perp}=0$. Since $y^{\parallel},f(y)^{\parallel}\in V$, we can write them as
@@
Now, let $\partial X$ be $C^{2}$-smooth. Notice that the tangent space of $Y$ at the point $x+f(x)$ for $x\in\partial X$ can be described as $\{\xi+\nabla_{\xi}f:\xi\in T_{x}\partial X\}$. Consider the symplectic subspaces $V=\mathop{\text{span}}\{x,f(x)\}$ and $V^{\omega}$. Remember that $f(x)$ and $x$ give characteristic directions in $T_{x}\partial X$ and $T_{f(x)}\partial X$ respectively. Therefore, by the definition of characteristic direction we have
@@
Every line of the form $\{x+tf(x):t\in\mathbb{R}\}$ intersects $Y$ exactly at two points $x+f(x)$ and xf(x)x-f(x) for every $x\in\partial X$. If $\partial X$ is $C^{2}$-smooth, then every such line is transversal to $Y$.
@@
Now, let $\partial X$ be $C^{2}$-smooth. Arguing by contradiction, assume that the line $\{x+tf(x):t\in\mathbb{R}\}$ intersects $Y$ not transversally at the point $x+f(x)$ for $x\in\partial X$ (case xf(x)x-f(x) is analogous). Following the proof of Theorem 4.2, there exists $\xi\in T_{x}\partial X$ such that
@@ 6. Discussion and open questions
In dimension two, for bodies with a $C^{2}$-smooth boundary, the answer is negative. Indeed, consider a symplectically self-polar convex body $X\subset\mathbb{R}^{2}$ and the corresponding set $Y$. Then $\partial X$ can be recovered from $Y$ by the so-called area construction (the outer billiard analog of the string construction [16]) as follows. Let $A$ be the area of the region bounded by $Y$. For any point $y$, consider the segment $y,z(y)],z(y)\in Y$ cutting from $Y$ a segment of a constant area equal to $\frac{A-4}{4}$ (Figure 2, gray area). Then,
@@
Figure 2. Area construction applied to $Y$ with the area parameter $\frac{A-4}{4}$ .
@@
In [7], analytic examples of Radon curves were found via a Lamé equation. Interestingly, are there analytic examples of symplectically self-polar bodies in higher dimensions?